# Tianxiang Chen USTC, China Alibaba Group, China

### 08 Mar 2024 v2

@@ 3.2 Mamba-in-Mamba for Efficient ISTD
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $4\times 4$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words. Unlike ViTs, we do not add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [28].
@@ 4.3 Visualization
##\##1 Visual Word $\rightarrow 2\times 2$ 78.71 76.43 99.12 5.965
##\##2 Visual Word $\rightarrow 8\times 8$ 79.50 77.68 100 2.980
@@
##\##5 Ours (VW $\rightarrow 4\times 4$, VS $\rightarrow H/8\times W/8$ ) 80.80 80.20 100 2.179
@@ 4.4 Ablation study
To explore the impact of information contained in each visual word, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $2\times 2$ (No.#$\#1$) or $8\times 8$ (No.#$\#2$) pixel region in the original image, in contrast to $4\times 4$ of our present setting. We also change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#3$) and $\frac{H}{16}\times\frac{W}{16}$ (No.#$\#4$) to ablate the effect of visual sentence granularity. We can observe that too large or too small division granularity cannot bring the most ideal performance. We adopt the configuration that demonstrates the best performance in our present setting.

### 17 Mar 2024 v3

# Tianxiang Chen USTC, China Alibaba Cloud, China

@@ 3.2 Mamba-in-Mamba for Efficient ISTD
We construct a convolutional stem, where a stack of $3\times 3$ convolutions is utilized, to produce visual words $\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$ and visual sentences $\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times D}$ at the first stage, where $C$ is the visual word dimension and $D$ is the visual sentence dimension. Each visual word corresponds to a $2\times 2$ pixel region in the original image, and each visual sentence is composed of $4\times 4$ visual words at this stage. Unlike ViTs, we do not add the position embedding bias to visual words and sentences due to the causal nature of visual mamba block [28].
@@ 4.3 Visualization
##\##1 Visual Word $\rightarrow 1\times 1$ 78.71 76.43 99.12 5.965
##\##2 Visual Word $\rightarrow 4\times 4$ 79.50 77.68 100 2.980
@@
##\##5 Ours (VW $\rightarrow 2\times 2$, VS $\rightarrow H/8\times W/8$ ) 80.80 80.20 100 2.179
@@ 4.4 Ablation study
To explore the impact of information contained in each visual word produced by the convolutional stem, we adjust the representation granularity of sub-patches so that each visual word corresponds to a $1\times 1$ (No.#$\#1$) or $4\times 4$ (No.#$\#2$) pixel region in the original image, in contrast to $2\times 2$ of our present setting. We also change the initial spatial shape of visual sentence from $\frac{H}{8}\times\frac{W}{8}$ to $\frac{H}{4}\times\frac{W}{4}$ (No.#$\#3$) and $\frac{H}{16}\times\frac{W}{16}$ (No.#$\#4$) to ablate the effect of visual sentence granularity. We can observe that too large or too small division granularity cannot bring the most ideal performance. We adopt the configuration that demonstrates the best performance in our present setting.